# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v3

@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable general-purpose task-solving capabilities across various fields and have been gradually incorporated into daily life. The development of LLMs is thriving, yet deploying LLMs on local devices presents practical challenges, including computational, memory, and energy resource limitations (Su et al., 2024; Wang et al., 2024). These issues hinder the deployment of large-scale LLMs on users’ local devices. Industry giants have developed strategies that either (i) involve lightweight versions of LLMs, such as Google’s Gemma and Meta’s LLaMA, which offer solutions with varying numbers of parameters (typically 3B, 8B, 13B, etc.) for fast and cost-efficient response, or (ii) place LLMs in the cloud, where users interact with them via the Internet, as seen with the ChatGPT application on mobile phones. The former often suffers from inferior inference performance, while the latter requires substantial network bandwidth and incurs high costs from service providers. This creates a pressing need to build systems that combine the advantages of both strategies, i.e., enhancing response quality while ensuring lower latency and usage costs.
@@ 1.1. Research Challenges
2. Understanding relationships between modalities, tasks, and dialogues: There are intricate associations between multi-modal data sources, inference tasks, and their temporal variation across multiple dialogues. Although various types of input data can improve the response quality of LLMs, not all tasks require comprehensive multi-modal data. For example, a message editing task may not require images as inputs, whereas some human activity tasks require various types of image modalities (Lee et al., 2024). Furthermore, when considering multi-dialogue scenarios, where each dialogue represents a task in a sequential order, earlier uploaded data modalities may provide sustained informational advantages, yet they might necessitate re-uploading if the information becomes outdated.
3. Uncertainty issues in offline training with LLM inference data: It is often desirable to train auxiliary models using LLM inference results, e.g., to learn modality-task associations. The computationally expensive nature of LLMs renders it impractical to conduct online training of such models. Consequently, generating datasets for offline training becomes necessary. However, due to inherent uncertainty in LLM inference, which arises from the probabilistic nature of language model outputs (Hu et al., 2023), such datasets may contain multiple identical samples with varying response scores. This uncertainty substantially complicates the evaluation of LLM outputs, particularly in multi-task or generative scenarios where standard answers or evaluation metrics do not exist.
@@ 2. Related Work
Efficient On-Device and Over-Network LLM. The rapid development of LLMs has profoundly influenced human-AI interaction while introducing significant challenges related to operational costs, energy consumption, and environmental impact. Researchers propose various solutions to enable deployment in resource-constrained devices (Chen et al., 2023), such as architectural optimization (Chu et al., 2023) and quantization (Lin et al., 2024). Furthermore, network-based approaches have emerged, such as (Shen et al., 2024) proposed the client-edge co-inference method, which partitions LLM into lightweight layers running on devices and parameter-heavy layers operating at the edge to enhance inference efficiency.
@@
However, existing frameworks primarily focus on either single-modal data sources or single-query accuracy (Zhang et al., 2024), neglecting the complex trade-off optimization of key resource indicators (e.g., response quality, latency, costs) (Dong et al., 2024) and failing to address the intricate relationships between multi-modal, multi-task, and multi-dialogue context (i.e., RQ1 and RQ2). Furthermore, (Zhang et al., 2024; He et al., 2024) employ datasets with fixed evaluation metrics, which may not fully represent the complexity in many real-world LLM applications. Moreover, online training methods incur prohibitively expensive costs and lead to irreproducible resource consumption, particularly when considering human subjects or alternative approaches for evaluating LLM inference (i.e., RQ3). In contrast to previous work, we propose an offline training approach to preserve LLM inference results. While this inevitably introduces uncertainties in the LLM inference data, we address this challenge by further incorporating an uncertainty estimator for response scores.
@@
Inference Selection and Offloading Optimization. More generally, there is a rich set of literature on strategies for optimizing local-edge-cloud inference location selection/offloading, based on metrics such as performance, cost, and latency (Singhal et al., 2024). These works explore diverse methodological approaches that can be categorized into network optimization strategies, such as confidence-based device-server selection (Al-Atat et al., 2024) and energy-delay considerations for LLM training (Liu and Zhao, 2024), as well as RL frameworks, such as accuracy-delay-fairness device-server selection (Beytur et al., 2024) and multi-device offloading with quality-latency-energy considerations (Chen et al., 2024).
@@
However, existing hierarchical inference approaches (Al-Atat et al., 2024; Beytur et al., 2024), which orchestrate offloading between local devices and cloud servers, only consider model size differences. They neglect fundamental distinctions in their input modalities, for example, one scenario is device-side LLMs being limited to text-only inputs, while server-side models can process multimodal inputs (i.e., RQ1). Furthermore, recall that the unique challenge in our settings is multi-dialogue interactions, yet these methods primarily focus on meeting specific performance standards or finding immediate optimal solutions without considering cumulative reward maximization throughout the decision-making process. (He et al., 2024) employed active inference for LLM inference offloading but disregarded multi-modal data and mutli-task in multi-dialogues, failing to align with practical LLM scenarios (i.e., RQ2). Differently from prior works, we implement these concepts in a multi-dialogue LLM scenario, not only maximizing cumulative rewards, but also considering the modality-task associations to select the optimal LLM and multi-modal data sources.
@@ 3.1. Multi-Modal, Multi-Task, Multi-Dialogue Interaction
During the conversation with LLMs, a user may ask additional questions using prompts potentially related to the previous dialogues (e.g., “Where did I put my dishes?” or “Please turn on all the lights in the kitchen”). Here, we assume each text prompt belongs to one of $N$ different task categories (e.g., assistant, recommendation, query, and message editing). Moreover, distinct tasks have varying levels of difficulty and require different types of multi-modal information to solve. For example, the prompt “Where did I put my dishes?” requires additional modalities (e.g., images) for the LLM to solve the task, while the prompt “Please turn on all the lights in the kitchen” may not require other sources of data, indicating that the task is simpler. These characteristics within a multi-task, multi-dialogue scenario with multi-modal data sources distinguish our setup from existing works (Zhang et al., 2024; He et al., 2024; Yang et al., 2024) that consider simpler settings (e.g., a single dialogue with text only), motivating the need for a new solution.
@@ 3.3.2. MDP and Optimization
In deep RL optimization, various strategies can be employed to adjust the parameters $\theta$ of the policy $\pi_{\theta}(a|s)$ with the goal of maximizing $J(\theta)$. To achieve this, a loss function $f(\theta)$ is defined such that minimizing $f(\theta)$ leads to maximizing $J(\theta)$. Taking Advantage Actor Critic (A2C) (Mnih et al., 2016) as an example, the loss function $f_{\text{A2C}}(\theta)$ includes both the policy loss and the value loss, along with an entropy term to encourage exploration:
@@
where $A(s,a)=r+\gamma V^{\pi_{\theta}}(s^{\prime})-V^{\pi_{\theta}}(s)$ for $s^{\prime}\neq s$ is an estimate of the advantage function, providing a measure of the relative value of taking action $a$ in state $s$. The coefficient $\zeta$ controls the weight of the entropy bonus $H(\pi_{\theta}(\cdot|s))$. In Sec. 4.2, we also consider Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Deep Q Network (DQN) (Mnih et al., 2015) versions of Eq. (8).
@@ 3.4. Response Score Uncertainty Estimation
Due to the vast number of parameters in LLMs, offline training approaches have become essential to mitigate substantial inference costs and enable future reusability (Ouyang et al., 2022). TMO’s RCRL procedure requires a specialized dataset for learning and exploration of LLM performance in multi-modal, multi-task, and multi-dialogue settings. Moreover, given the nature of multi-task settings, particularly in generative tasks, there often exist no explicit evaluation metrics, such as accuracy, task success rate, or error rate. Furthermore, the exploratory nature of RL, which frequently encounters new state-action pairs, presents significant evaluation challenges.
@@ 3.4.1. Response Quality Evaluation
The response score $S_{R_{t}}$ measures the response quality generated by the selected LLM in each dialogue. In multi-task scenarios, evaluating the response quality requires assessing the actual utility of responses in aiding human users, particularly in query tasks. We employ the quantitative evaluation scheme outlined by (Liu et al., 2024), illustrated on the right side of Fig. 3. This involves using two additional text-only LLMs specifically to assess the quality of candidate LLM services:
@@ 3.4.2. Handling Uncertainty in Response Score Estimation
In contrast to (Ran et al., 2023), which aims to avoid selecting OOD actions, our goal is to more accurately estimate the response quality under OOD actions rather than constrain them. This adds flexibility for datasets where the states and actions are randomly generated (e.g., the current task, selected LLM, and selected modality), making it inevitable that some state-action pairs will not appear in the dataset (such as for our M4A1 dataset described in Sec. 4.1).
@@ 3.5. Task-Modality Association
TMO is designed to operate in multi-modal, multi-task scenarios. In such settings, quantifying the relevance between users’ text prompts and multi-modal data sources enables selection of the most pertinent multi-modal data sources to offload to the cloud LLM. Consequently, we employ feature extractors to transform both tasks and modalities into feature vectors of identical dimensionality for similarity computation. Given this premise, we leverage the pre-trained CLIP model (Radford et al., 2021), a transformer-based architecture with text and image encoders, to extract features from prompts and multi-model data sources. We employ a normalized cross-modality similarity metric to calculate the association score between a text prompt $P$ and data modality $m$, which is defined as:
@@ 3.6.1. Local Computation Latency $\psi_{\text{Local}}$
The computational latency of the local LLM, or say inference time, is influenced not only by the number of LLM parameters, but also significantly by the length of the prompt, denoted $|P|$, and of the response, $|R|$. Following (Canziani et al., 2016), we assume that the computational latency is proportional to the number of parameters. If the local LLM is comprised of $|\text{LLM}_{\text{Local}}|$ parameters and requires $2|\text{LLM}_{\text{Local}}|$ floating point operations (FLOPS) for forward propagation, the computational demand for each token in a reference processing length $|P|_{\text{ref}}$ is calculated as $\frac{2|\text{LLM}_{\text{Local}}|}{|P|_{\text{ref}}}$ FLOPS. The latency calculation considers the theoretical peak performance of the local device’s GPU, $TF_{\text{peak}}$ FLOPS. The computational latency can be computed as:
@@ 3.7.2. Cloud LLM Usage Function $\phi_{\text{Cloud}}$
For cloud-based LLMs, in addition to the inference costs of prompts and responses, the inference costs of data modalities must also be considered. These costs are typically available on the service provider’s website. Taking GPT-4o as an example, the prompt cost rate is $\varphi_{P}=\$0.005/1K$ tokens, the response cost rate is $\varphi_{R}=\$0.015/1K$ tokens, and the data modality costs are proportional to their size and resolution at a rate $\varphi_{m}$ for modality $m\in\mathcal{M}$. Therefore, the usage cost of the cloud LLM can be defined as follows:
@@ 4.2. Experiment Setup
Training Setup. The M4A1 dataset is split into a training set and a test set in an 8:2 ratio. The training set is used to train RLs while the test set is used to evaluate the performance at inference. When implementing our approach, we consider several popular discrete RL algorithms, including PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), and A2C (Mnih et al., 2016). We consider these methods with and without resource constraints (corresponding to cases with $\lambda>0$ and $\lambda=0$ in Eq. (10)), and denote RC-PPO, RC-DQN, and RC-A2C as the methods with the resource constraints. The policies are parameterized by multilayer perceptrons (MLP), with a total of 30,000 time steps set for training. For each experiment, we repeat the process three times and report the standard deviation as the error bar. We implement the TMO system and the baselines in PyTorch and conducted the experiments on an NVIDIA A100 GPU with 40 GB of memory.
Baselines. We compare the RL policies with three naive baselines: random (with random LLM and modality selection), local (using only the local LLM), and cloud (using only the cloud LLM, but with random modality selection). Inspired by recent advances in LLM-as-Agent (Xi et al., 2023; Liu et al., 2023), we substitute the RL agent with an LLM agent as TMO’s decision engine for the selection of LLMs and modalities. Specifically, current states and possible actions are translated into natural language descriptions as contextual input for the LLM agents, which then output an action in response to this prompt, following an RL-like procedure. The LLM agents include Phi-3-mini, Phi-3.5-mini, LLaMA-3.2-3B, LLaMA-3.1-8B, Mistral-7B-v0.3, FLAN-T5-large, FLAN-T5-xl, Gemma-2-2B, Gemma-2-9B, GPT-3.5-turbo, GPT-4o-mini, GPT-4o, OpenAI o1-mini, OpenAI o1, and OpenAI o3-mini. We reiterate that none of the existing works can be directly applied to our setting, which features multi-modal, multi-task, and multi-dialogue characteristics. Therefore, we reimplement AIwRG (He et al., 2024), an active inference method, and PerLLM (Yang et al., 2024), a bandit-based approach using upper confidence bound (UCB) for exploration-exploitation trade-off, integrating the settings introduced in this paper to adapt them to our research problem.
@@
LLM-as-Agent (Xi et al., 2023; Liu et al., 2023) for Inference Offloading (Ignored LLM Agent’s Latency and Usage Cost)
@@
AIwRG (He et al., 2024) 1.02 $\pm$0.05 23.73 $\pm$0.98 44.01 $\pm$3.50 0.69 $\pm$0.06 113 244 0 0 3852 4.68 $\pm$0.18 9.26 $\pm$1.83
PerLLM (Yang et al., 2024) 0.97 $\pm$0.06 21.12 $\pm$0.08 1.60 $\pm$0.01 0.66 $\pm$0.06 281 3922 0 0 0 0.00 $\pm$0.00 0.00 $\pm$0.00

### v4

@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable general-purpose task-solving capabilities across various fields and have been gradually incorporated into daily life. The development of LLMs is thriving, yet deploying LLMs on local devices presents practical challenges, including computational, memory, and energy resource limitations (su2024titanic; wang2024resource). These issues hinder the deployment of large-scale LLMs on users’ local devices. Industry giants have developed strategies that either (i) involve lightweight versions of LLMs, such as Google’s Gemma and Meta’s LLaMA, which offer solutions with varying numbers of parameters (typically 3B, 8B, 13B, etc.) for fast and cost-efficient response, or (ii) place LLMs in the cloud, where users interact with them via the Internet, as seen with the ChatGPT application on mobile phones. The former often suffers from inferior inference performance, while the latter requires substantial network bandwidth and incurs high costs from service providers. This creates a pressing need to build systems that combine the advantages of both strategies, i.e., enhancing response quality while ensuring lower latency and usage costs.
@@ 1.1. Research Challenges
2. Understanding relationships between modalities, tasks, and dialogues: There are intricate associations between multi-modal data sources, inference tasks, and their temporal variation across multiple dialogues. Although various types of input data can improve the response quality of LLMs, not all tasks require comprehensive multi-modal data. For example, a message editing task may not require images as inputs, whereas some human activity tasks require various types of image modalities (lee2024gazepointar). Furthermore, when considering multi-dialogue scenarios, where each dialogue represents a task in a sequential order, earlier uploaded data modalities may provide sustained informational advantages, yet they might necessitate re-uploading if the information becomes outdated.
3. Uncertainty issues in offline training with LLM inference data: It is often desirable to train auxiliary models using LLM inference results, e.g., to learn modality-task associations. The computationally expensive nature of LLMs renders it impractical to conduct online training of such models. Consequently, generating datasets for offline training becomes necessary. However, due to inherent uncertainty in LLM inference, which arises from the probabilistic nature of language model outputs (hu2023uncertainty), such datasets may contain multiple identical samples with varying response scores. This uncertainty substantially complicates the evaluation of LLM outputs, particularly in multi-task or generative scenarios where standard answers or evaluation metrics do not exist.
@@ 2. Related Work
Efficient On-Device and Over-Network LLM. The rapid development of LLMs has profoundly influenced human-AI interaction while introducing significant challenges related to operational costs, energy consumption, and environmental impact. Researchers propose various solutions to enable deployment in resource-constrained devices (chen2023frugalgpt), such as architectural optimization (chu2023mobilevlm) and quantization (lin2024awq). Furthermore, network-based approaches have emerged, such as (shen2024large) proposed the client-edge co-inference method, which partitions LLM into lightweight layers running on devices and parameter-heavy layers operating at the edge to enhance inference efficiency.
@@
However, existing frameworks primarily focus on either single-modal data sources or single-query accuracy (zhang2024treacle), neglecting the complex trade-off optimization of key resource indicators (e.g., response quality, latency, costs) (dong2024creating) and failing to address the intricate relationships between multi-modal, multi-task, and multi-dialogue context (i.e., RQ1 and RQ2). Furthermore, (zhang2024treacle; he2024large) employ datasets with fixed evaluation metrics, which may not fully represent the complexity in many real-world LLM applications. Moreover, online training methods incur prohibitively expensive costs and lead to irreproducible resource consumption, particularly when considering human subjects or alternative approaches for evaluating LLM inference (i.e., RQ3). In contrast to previous work, we propose an offline training approach to preserve LLM inference results. While this inevitably introduces uncertainties in the LLM inference data, we address this challenge by further incorporating an uncertainty estimator for response scores.
@@
Inference Selection and Offloading Optimization. More generally, there is a rich set of literature on strategies for optimizing local-edge-cloud inference location selection/offloading, based on metrics such as performance, cost, and latency (singhal2024resource). These works explore diverse methodological approaches that can be categorized into network optimization strategies, such as confidence-based device-server selection (al2024regret) and energy-delay considerations for LLM training (liu2024resource), as well as RL frameworks, such as accuracy-delay-fairness device-server selection (beytur2024optimization) and multi-device offloading with quality-latency-energy considerations (chen2024tilesr).
@@
However, existing hierarchical inference approaches (al2024regret; beytur2024optimization), which orchestrate offloading between local devices and cloud servers, only consider model size differences. They neglect fundamental distinctions in their input modalities, for example, one scenario is device-side LLMs being limited to text-only inputs, while server-side models can process multimodal inputs (i.e., RQ1). Furthermore, recall that the unique challenge in our settings is multi-dialogue interactions, yet these methods primarily focus on meeting specific performance standards or finding immediate optimal solutions without considering cumulative reward maximization throughout the decision-making process. (he2024large) employed active inference for LLM inference offloading but disregarded multi-modal data and mutli-task in multi-dialogues, failing to align with practical LLM scenarios (i.e., RQ2). Differently from prior works, we implement these concepts in a multi-dialogue LLM scenario, not only maximizing cumulative rewards, but also considering the modality-task associations to select the optimal LLM and multi-modal data sources.
@@ 3.1. Multi-Modal, Multi-Task, Multi-Dialogue Interaction
During the conversation with LLMs, a user may ask additional questions using prompts potentially related to the previous dialogues (e.g., “Where did I put my dishes?” or “Please turn on all the lights in the kitchen”). Here, we assume each text prompt belongs to one of $N$ different task categories (e.g., assistant, recommendation, query, and message editing). Moreover, distinct tasks have varying levels of difficulty and require different types of multi-modal information to solve. For example, the prompt “Where did I put my dishes?” requires additional modalities (e.g., images) for the LLM to solve the task, while the prompt “Please turn on all the lights in the kitchen” may not require other sources of data, indicating that the task is simpler. These characteristics within a multi-task, multi-dialogue scenario with multi-modal data sources distinguish our setup from existing works (zhang2024treacle; he2024large; yang2024perllm) that consider simpler settings (e.g., a single dialogue with text only), motivating the need for a new solution.
@@ 3.3.2. MDP and Optimization
In deep RL optimization, various strategies can be employed to adjust the parameters $\theta$ of the policy $\pi_{\theta}(a|s)$ with the goal of maximizing $J(\theta)$. To achieve this, a loss function $f(\theta)$ is defined such that minimizing $f(\theta)$ leads to maximizing $J(\theta)$. Taking Advantage Actor Critic (A2C) (mnih2016asynchronous) as an example, the loss function $f_{\text{A2C}}(\theta)$ includes both the policy loss and the value loss, along with an entropy term to encourage exploration:
@@
where $A(s,a)=r+\gamma V^{\pi_{\theta}}(s^{\prime})-V^{\pi_{\theta}}(s)$ for $s^{\prime}\neq s$ is an estimate of the advantage function, providing a measure of the relative value of taking action $a$ in state $s$. The coefficient $\zeta$ controls the weight of the entropy bonus $H(\pi_{\theta}(\cdot|s))$. In Sec. 4.2, we also consider Proximal Policy Optimization (PPO) (schulman2017proximal) and Deep Q Network (DQN) (mnih2015human) versions of Eq. (8).
@@ 3.4. Response Score Uncertainty Estimation
Due to the vast number of parameters in LLMs, offline training approaches have become essential to mitigate substantial inference costs and enable future reusability (ouyang2022training). TMO’s RCRL procedure requires a specialized dataset for learning and exploration of LLM performance in multi-modal, multi-task, and multi-dialogue settings. Moreover, given the nature of multi-task settings, particularly in generative tasks, there often exist no explicit evaluation metrics, such as accuracy, task success rate, or error rate. Furthermore, the exploratory nature of RL, which frequently encounters new state-action pairs, presents significant evaluation challenges.
@@ 3.4.1. Response Quality Evaluation
The response score $S_{R_{t}}$ measures the response quality generated by the selected LLM in each dialogue. In multi-task scenarios, evaluating the response quality requires assessing the actual utility of responses in aiding human users, particularly in query tasks. We employ the quantitative evaluation scheme outlined by (liu2024visual), illustrated on the right side of Fig. 3. This involves using two additional text-only LLMs specifically to assess the quality of candidate LLM services:
@@ 3.4.2. Handling Uncertainty in Response Score Estimation
In contrast to (ran2023policy), which aims to avoid selecting OOD actions, our goal is to more accurately estimate the response quality under OOD actions rather than constrain them. This adds flexibility for datasets where the states and actions are randomly generated (e.g., the current task, selected LLM, and selected modality), making it inevitable that some state-action pairs will not appear in the dataset (such as for our M4A1 dataset described in Sec. 4.1).
@@ 3.5. Task-Modality Association
TMO is designed to operate in multi-modal, multi-task scenarios. In such settings, quantifying the relevance between users’ text prompts and multi-modal data sources enables selection of the most pertinent multi-modal data sources to offload to the cloud LLM. Consequently, we employ feature extractors to transform both tasks and modalities into feature vectors of identical dimensionality for similarity computation. Given this premise, we leverage the pre-trained CLIP model (radford2021learning), a transformer-based architecture with text and image encoders, to extract features from prompts and multi-model data sources. We employ a normalized cross-modality similarity metric to calculate the association score between a text prompt $P$ and data modality $m$, which is defined as:
@@ 3.6.1. Local Computation Latency $\psi_{\text{Local}}$
The computational latency of the local LLM, or say inference time, is influenced not only by the number of LLM parameters, but also significantly by the length of the prompt, denoted $|P|$, and of the response, $|R|$. Following (canziani2016analysis), we assume that the computational latency is proportional to the number of parameters. If the local LLM is comprised of $|\text{LLM}_{\text{Local}}|$ parameters and requires $2|\text{LLM}_{\text{Local}}|$ floating point operations (FLOPS) for forward propagation, the computational demand for each token in a reference processing length $|P|_{\text{ref}}$ is calculated as $\frac{2|\text{LLM}_{\text{Local}}|}{|P|_{\text{ref}}}$ FLOPS. The latency calculation considers the theoretical peak performance of the local device’s GPU, $TF_{\text{peak}}$ FLOPS. The computational latency can be computed as:
@@ 3.7.2. Cloud LLM Usage Function $\phi_{\text{Cloud}}$
For cloud-based LLMs, in addition to the inference costs of prompts and responses, the inference costs of data modalities must also be considered. These costs are typically available on the service provider’s website. Taking GPT-4o as an example, the prompt cost rate is $\varphi_{P}=\mathdollar 0.005/1K$ tokens, the response cost rate is $\varphi_{R}=\mathdollar 0.015/1K$ tokens, and the data modality costs are proportional to their size and resolution at a rate $\varphi_{m}$ for modality $m\in\mathcal{M}$. Therefore, the usage cost of the cloud LLM can be defined as follows:
@@ 4.2. Experiment Setup
Training Setup. The M4A1 dataset is split into a training set and a test set in an 8:2 ratio. The training set is used to train RLs while the test set is used to evaluate the performance at inference. When implementing our approach, we consider several popular discrete RL algorithms, including PPO (schulman2017proximal), DQN (mnih2015human), and A2C (mnih2016asynchronous). We consider these methods with and without resource constraints (corresponding to cases with $\lambda>0$ and $\lambda=0$ in Eq. (10)), and denote RC-PPO, RC-DQN, and RC-A2C as the methods with the resource constraints. The policies are parameterized by multilayer perceptrons (MLP), with a total of 30,000 time steps set for training. For each experiment, we repeat the process three times and report the standard deviation as the error bar. We implement the TMO system and the baselines in PyTorch and conducted the experiments on an NVIDIA A100 GPU with 40 GB of memory.
Baselines. We compare the RL policies with three naive baselines: random (with random LLM and modality selection), local (using only the local LLM), and cloud (using only the cloud LLM, but with random modality selection). Inspired by recent advances in LLM-as-Agent (xi2023rise; liu2023agentbench), we substitute the RL agent with an LLM agent as TMO’s decision engine for the selection of LLMs and modalities. Specifically, current states and possible actions are translated into natural language descriptions as contextual input for the LLM agents, which then output an action in response to this prompt, following an RL-like procedure. The LLM agents include Phi-3-mini, Phi-3.5-mini, LLaMA-3.2-3B, LLaMA-3.1-8B, Mistral-7B-v0.3, FLAN-T5-large, FLAN-T5-xl, Gemma-2-2B, Gemma-2-9B, GPT-3.5-turbo, GPT-4o-mini, GPT-4o, OpenAI o1-mini, OpenAI o1, and OpenAI o3-mini. We reiterate that none of the existing works can be directly applied to our setting, which features multi-modal, multi-task, and multi-dialogue characteristics. Therefore, we reimplement AIwRG (he2024large), an active inference method, and PerLLM (yang2024perllm), a bandit-based approach using upper confidence bound (UCB) for exploration-exploitation trade-off, integrating the settings introduced in this paper to adapt them to our research problem.
@@
LLM-as-Agent (xi2023rise; liu2023agentbench) for Inference Offloading (Ignored LLM Agent’s Latency and Usage Cost)
@@
AIwRG (he2024large) 1.02 $\pm$0.05 23.73 $\pm$0.98 44.01 $\pm$3.50 0.69 $\pm$0.06 113 244 0 0 3852 4.68 $\pm$0.18 9.26 $\pm$1.83
PerLLM (yang2024perllm) 0.97 $\pm$0.06 21.12 $\pm$0.08 1.60 $\pm$0.01 0.66 $\pm$0.06 281 3922 0 0 0 0.00 $\pm$0.00 0.00 $\pm$0.00